73. Consider the reaction:

CO(g) + 2H2(g) <--> CH3OH(g) Kp= 2.26X10E4 at 25 C

Calculate ΔGorxnfor the reaction at 25C under each of the following conditions:

a. standard conditions

b. at equilibrium

c. PCH3OH= 1.0 atm; PCO= PH2 = 0.010 atm
Answer #1

a. orxn = - RTlnKp [ T = 25oc = 298K]

= -8.314*298*ln(2.26*104)

= - 2477.57*10.02

= - 24839.4 J = - 24.839KJ

b. At equilibrium, orxn = 0. ( kp​​​​​​= 1)

C.

For the given reaction Kp = P(CH3OH)/ P(CO) *P2(H2)

= 1/(0.010)*(0.010)2

= 106

Then, orxn = - RT lnKp

= - 8.314*298ln(106)

= - 2477.57*13.81

= -34228.9 J

= - 34.22 KJ